# Aegis: A Decentralized Expansion Blockchain

### v2

@@ 2.4 Chain Combining
Fitzi et al. [27] and Wang et al. [61] combine multiple ledgers to create a ledger with better properties. However, unlike Aegis, all steps are taken within the ledgers, unlike Aegis, whose expansion chain can progress by expansion-chain consensus steps that are not immediately checkpointed.
2.5 Layer-2 Blockchains
Various systems utilize a trusted primary blockchain to validate and secure second-level applications. This includes rollups that increase the primary chain’s throughput, and offer lower latency and cost using optimistic [31, 3, 47] and zero-knowledge [56, 39] methodologies. Gudgeon et al. [30] provide a comprehensive systemization of knowledge. However, while the validity of the operations of those systems is guaranteed by the underlying blockchain, their content is, by and large, approved by centralized services [39, 47]. In contrast to those solutions, Aegis implements a decentralized sequencing mechanism, which can be used by any of those rollups together with their on-chain validation mechanisms.
Other solutions include so-called sidechains and payment channels. Sidechains (e.g., [48, 29]) are independent blockchains with mechanisms to transact with a primary chain, but their security is independent of the primary chain, unlike the expansion chain in Aegis. Payment channels are point-to-point channels [49, 24], which can aggregate several payments between two parties. These can also be used to form a payment network, allowing payments between any two parties in the network, without the need for a direct channel between them. But, as each channel is exclusively controlled by its two counterparties, payment channels are not decentralized, like Aegis.
3 Model
@@
The system (§3.1) includes a primary blockchain and a set of nodes. We assume a generic consensus algorithm (§3.2) that the nodes use aiming to form a secondary blockchain (§3.3). Appendix A summarizes the notation.
3.1 Principals and Network
@@
The system comprises an unbounded set of nodes (as in [34, 50]) and a primary blockchain (or simply primary) denoted by $\underline{\mathcal{L}}$. Time progresses in steps. In each step a node receives messages from previous steps, executes local computations, updates its state, and sends messages.
@@
The primary chain is implemented by a trusted party that maintains a state including token balance for each node and a so-called staked amount for each node. Initially, each node has some initial token balance and zero staked tokens. The primary also maintains node-defined smart contracts implementing arbitrary automatons. Nodes interact with the primary by issuing transactions that update its state and interact with smart contracts. The primary chain aggregates transactions into blocks and extends the blockchain at set intervals. Its state is the result of parsing the series of all transactions. A node cannot guess the hash of a future block except with negligible probability.
@@
A node can issue a transaction to stake, i.e., deposit some of their tokens, or unstake, i.e., order the withdrawal of tokens she had previously staked. Unstaking completes $\Delta_{\textit{active}}$ steps after the order transaction is placed in the chain. While a node’s tokens are staked, the node is considered active. During this period, a node can either follow the protocol (and be considered correct) or misbehave arbitrarily. Correct nodes adhere to the protocol while they are staked, if misbehavior could result in losing their stake. However, once they unstake, they are no longer bound by the protocol and can behave arbitrarily without risk of losing their stake.
@@
The set of nodes that have staked but not yet unstaked after a specific block is called a committee; until the time any of them completed unstaking, the committee is active. In all committees, the ratio of stake held by correct nodes is assumed to be greater than a constant threshold, denoted by $\alpha$.
@@
Communication among nodes is reliable and partially synchronous: There exists a time $t_{\textit{GST}}$, unknown to the nodes, and message delivery is only bounded after it. That is, if a message is sent at time $t$ then it arrives at all nodes by time ${\max(t,t_{\textit{GST}})+\Delta_{\textit{prop}}}$. In particular, blocks published by the nodes are delivered to all nodes, even those that stake much later. [^3] In practice [10], this network is implemented by the nodes’ peer-to-peer network, where nodes keep all published blocks available.
After $t_{\textit{GST}}$ the set of active nodes stabilizes, no one stakes or unstakes.
@@
Communication with the primary chain is synchronous: All nodes can observe the current state of the primary chain and issue transactions that are added to it within a bounded time $\Delta_{\underline{\mathcal{L}}\textit{-write}}$. This assumption is stronger than the one for internode communication, but it is more reasonable in this case because we rely on an already established primary chain. In practice, this assumption is met by choosing conservative bounds for the time it takes to write to the primary chain.
3.2 Consensus Algorithm
@@
We assume we are given an asynchronous (single-instance) Byzantine fault tolerant consensus protocol (e.g., [12, 15]) implemented in a function consensusStep. Called by a process, consensusStep takes steps in the protocol, maintaining a local state as necessary. It is executed by a weighted set of nodes $\mathcal{N}$, given as a parameter. (We will later use the staked amount as the weight, thus $\mathcal{N}$ will constitute a committee.) It takes a consensus instance ID, which is a pair of the set of nodes $\mathcal{N}$ and a nonce id to distinguish different instances. It also takes a block $b$ the caller wishes to propose. It returns either a block if it reached a decision or $\bot$ otherwise: $\textit{consensusStep}(b;\mathcal{N},\textit{id})\rightarrow\textit{block}\textup{ or }\bot$.
@@
Validity
If all nodes in $\mathcal{N}$ are correct and they all execute consensusStep with the same ID and the same proposal $b$, then their consensusStep call never returns a $\bot$ value $b^{\prime}\neq b$.
Termination
@@
Finally, the protocol supports forensics, that is, it allows identifying misbehavior of the nodes [54, 20]. Specifically, if the number of Byzantine nodes is larger than the bound, they can cause an agreement violation. In such a case, the correct nodes can publish data allowing them to identify misbehaving parties.
3.3 Goal
The goal of the system is to form an expansion blockchain, denoted by $\mathcal{L}$. Each node $i$ has an input value $v_{i}$ for each position in the ledger. We do not consider the content of this blockchain, which could serve as a decentralized sequencer, oracle, etc. (see §2). We only consider its blocks, which are maintained by the nodes as local vectors. If a node stores a block in a position $k$ in its local vector we say it logs this block. Note that this is a chain, not a single-instance consensus as in Section 3.2. The expansion blockchain should achieve the following (referring to expansion-chain blocks and positions):
Agreement
@@ 3.3 Goal
Validity
If all nodes are correct and their inputs for all positions ${v_{1},v_{2},\dots}$ are the same, then in all positions $k$ no node logs a block with a different value $u_{k}\neq v_{k}$.
Progress
There exists a time $t>t_{\textit{GST}}$ such that after $t$, each correct node logs new blocks at a rate of at least $1/(\Delta_{\textit{consensus}}+\Delta_{\textit{prop}})$.
@@
4 Aegis
@@
We overview the protocol design (§4.1), then present the Aegis blockchain structure (§4.2) and the primary-chain contract (§4.3). We then describe the peer-to-peer protocol Aegis nodes execute to form the blockchain (§4.4), and discuss the execution of the forensics procedure to identify and penalize misbehaving nodes (§4.5).
4.1 Protocol Outline
Aegis includes two components: A protocol that Aegis nodes execute and a smart contract that the primary chain does. Roughly, the protocol executes a series of consensus instances logging blocks consecutively. Each block $b$ contains a reference (hash) to its parent. The blocks thus form a tree, rooted at an agreed-upon Genesis block, with each having a height $\textit{height}(b)$: the number of edges on the path from $b$ to the Genesis block.
Each block also contains a reference to a primary block $B$. The primary block implicitly specifies a committee—the set of nodes with staked tokens at the state of block $B$, weighted by the staked amounts. (Nodes must include contact information in the staking transaction, so other nodes would know where to send messages; we omit this detail to simplify the presentation.) This committee is, in the good case, the committee that chooses the block following $b$. For example, in Figure 1, Aegis block $A2$ references primary block $P4$, so the committee defined by block $P4$ chooses the subsequent Aegis block $A3$.
Occasionally, Aegis nodes checkpoint blocks by sending them to the Aegis smart contract on the primary chain. The contract receives the checkpointed block and accepts the checkpoint only if its committee is active.
If Aegis could not produce a block for a duration of $\Delta_{\textit{active}}$—that is, in the bad case—the previous committee might not be active. In this situation, a new committee must be selected to extend the chain. For this, some node issues a reset in the primary-chain Aegis contract. The contract would accept the reset only if there were no checkpoints or resets in the last $\Delta_{\textit{active}}$ steps. This initiates a new committee, specified by the block in which the reset occurs. This committee can produce a new block following the previous checkpointed Aegis block using the committee specified by the reset.
We proceed to describe the algorithm in detail.
4.2 Chain Structure

### v3

@@ 2. Related Work
Several previous works combine multiple ledgers to create a ledger with better properties (Fitzi et al., 2020; Wang et al., 2022; Tas et al., 2024). However, all steps are taken within the used ledgers, unlike Aegis, whose tethered chain can progress by tethered-chain consensus steps faster than the rate it is checkpointed.
3. Model
@@
The system (§3.1) includes a primary blockchain and a set of nodes. We assume a generic consensus algorithm (§3.2) that the nodes use aiming to form a tethered blockchain (§3.3). Appendix A summarizes the notation.
3.1. Principals and Network
@@
The system comprises an unbounded set of nodes (as in (Kiayias et al., 2017; Pu et al., 2023)) and a primary blockchain (or simply primary) denoted by $\underline{\mathcal{L}}$. Time progresses in steps. In each step a node receives messages from previous steps, executes local computations, updates its state, and sends messages.
@@
The primary chain is implemented by a trusted party that maintains a state including token balance for each node and a so-called staked amount for each node. Initially, each node has some initial token balance. The primary also maintains node-defined smart contracts implementing arbitrary automatons. Nodes interact with the primary by issuing transactions that update its state and interact with smart contracts. The primary chain aggregates transactions into blocks and extends the blockchain at set intervals. Its state is the result of parsing the series of all transactions. A node cannot guess the hash of a future block except with negligible probability.
@@
A node can issue a transaction to stake, deposit some of their tokens, or unstake, order the withdrawal of tokens she had previously staked. Unstaking completes $\Delta_{\textit{active}}$ steps after the order transaction is placed in the chain. While a node’s tokens are staked, the node is considered active. During this period, a node can either follow the protocol and be considered correct, or misbehave arbitrarily. Correct nodes adhere to the protocol while they are staked, if misbehavior could result in losing their stake. However, once they unstake, they are no longer bound by the protocol and can behave arbitrarily without the risk of losing their stake.
@@
The set of nodes that have staked but not yet unstaked after a specific block is called a committee; until the time any of them completed unstaking, the committee is active. In all committees, the ratio of stake held by correct nodes is greater than a threshold, denoted by $\alpha$.
@@
Communication among nodes is reliable and partially synchronous: In every execution there exists a time $t_{\textit{GST}}$, unknown to the nodes, and message delivery is bounded only after it. That is, if a message is sent at time $t$ then it arrives at all nodes by time ${\max(t,t_{\textit{GST}})+\Delta_{\textit{prop}}}$. In particular, blocks published by the nodes are delivered to all nodes, even those that stake much later [^3] In practice (Bitcoin community, 2013), this is implemented by the nodes’ peer-to-peer network, where nodes keep all published blocks available. .
@@
Communication with the primary chain is synchronous: All nodes can observe the current state of the primary chain and issue transactions that are added to it within a bounded time $\Delta_{\underline{\mathcal{L}}\textit{-write}}$. This assumption is stronger than the one for internode communication, but it is more reasonable in this case because we rely on an already established primary chain. In practice, this assumption is met by choosing conservative bounds for the time it takes to write to the primary chain, even in times of congestion in the primary chain and perhaps even in the face of a DoS attack.
System behavior is orchestrated by a scheduler. The scheduler determines message arrival times, subject to the constraints above. It also determines when nodes stake and unstake–to represent the fact this decision is exogenous to the protocol that the participating nodes execute.
3.2. Consensus Algorithm
@@
We are given a (single-instance) Byzantine fault tolerant consensus protocol designed for partial synchrony (e.g., (Dwork et al., 1988; Castro et al., 1999)) implemented in a function consensusStep. Called by a process, consensusStep takes steps in the protocol, maintaining a local state as necessary. It is executed by a weighted set of nodes $\mathcal{N}$, given as a parameter. (We will later use the staked amount as the weight, thus $\mathcal{N}$ will constitute a committee.) It takes a consensus instance ID, which is a pair of the set of nodes $\mathcal{N}$ and a nonce id to distinguish different instances. It also takes a block $b$ the caller wishes to propose. It returns either a block if it reached a decision or $\bot$ otherwise: $\textit{consensusStep}(b;\mathcal{N},\textit{id})\rightarrow\textit{block}\textup{ or }\bot$.
@@
Validity:
If all nodes in $\mathcal{N}$ are correct, and they all execute consensusStep with the same ID and the same proposal $b$, then their consensusStep call never returns a $\bot$ value $b^{\prime}\neq b$.
Termination:
@@
Finally, the protocol supports forensics, that is, it allows identifying misbehavior of the nodes (Sheng et al., 2021; Civit et al., 2021). Specifically, if the number of Byzantine nodes is larger than the bound, they can cause an agreement violation. In such a case, the correct nodes can publish data allowing them to identify misbehaving parties. To allow for enough time to run forensics before nodes in an active committee can unstake, we require that ${\Delta_{\textit{active}}>3\Delta_{\underline{\mathcal{L}}\textit{-write}}}$.
3.3. Goal
The goal of the system is to form a tethered blockchain, denoted by $\mathcal{L}$. Each node $i$ has a time-varying input value for each position of $\mathcal{L}$. For a height $k$ (distance from the root) and time $t$, it is denoted by $v^{i}_{t}(k)$. We do not consider the content of $\mathcal{L}$, which could serve as a decentralized sequencer, oracle, etc. (see §2). We only consider its blocks, which are maintained by the nodes as local vectors. If a node stores a block in a position $k$ in its local vector we say it logs this block. Note that this is a chain, not a single-instance consensus as in Section 3.2. The tethered blockchain should achieve the following 3 properties (referring to tethered-chain blocks and heights).
We use the standard Agreement property.
Agreement:
@@ 3.3. Goal
Agreement implies instant finality: Once a value is logged, it will not be overturned.
For validity, we require each decision to be based on fresh inputs. This embodies the interactive nature of the protocol, where each block is based on the previous one. It also prevents batching outside block limits, that is, the nodes cannot agree in a single consensus iteration on multiple blocks.
No-Batching validity:
Let $t^{\prime}$ be the time of the first decision at height $k$ (i.e., at least one node decides at height $k$ at time $t^{\prime}$, and no node decides at height $k$ before $t^{\prime}$). If (1) all input functions for height $k+1$ return value $v$ after $t^{\prime}$ (i.e., $\forall i,t^{\prime\prime}>t^{\prime}:v^{i}_{t^{\prime\prime}}(k+1)=v$); (2) all nodes are correct; and (3) there are no configuration changes; then no node decides on a value different from $v$ at height $k+1$.
Finally, we require rapid progress after the protocol stabilizes at some time after $t_{\textit{GST}}$.
Progress:
There exists a protocol stabilization bound $\Delta_{\textit{stabilize}}$ such that for all executions, for all ${t>t_{\textit{GST}}+\Delta_{\textit{stabilize}}}$, if the latest decided value by any node at $t$ is $k$, then by $t+\Delta_{\textit{consensus}}+\Delta_{\textit{prop}}$ all correct nodes decide on ${k+1}$
@@
4. Bounded Primary-Chain Access is Necessary
We now show that Aegis is optimal in the following sense: There is no solution to the problem without synchronous access to the primary ledger. Formally, assume for the rest of this section that instead of synchronous access to the primary ledger, there are no guarantees before $t_{\textit{GST}}$. That is, a write order to the primary at time $t$ is only guaranteed to be completed by time ${\max\{t,t_{\textit{GST}}\}+\Delta_{\underline{\mathcal{L}}\textit{-write}}}$. Note that all messages ever sent and all primary-writes ever completed are visible to newly joining nodes.
In addition, for the sake of simplicity, in the rest of the section, we assume that solutions can only log one bit, either  0 or 1. Our results, trivially transfer to the case where more values are possible.
4.1. Naive Attempts
For grasping the problem at hand, we suggest a few naive attempts which fail to achieve the desired properties.
One possible approach for a solution is for each node to submit to the primary ledger a proposal for each height with incrementing order. The first value at each height is the new bit, so all nodes log the same values. Only this extends at most at a rate of $\Delta_{\underline{\mathcal{L}}\textit{-write}}$ even after $t_{\textit{GST}}$, since this is the time (bound) to find the block at each height (the time to write a block to the ledger); and doesn’t satisfy the Progress property.
Another possibility is for nodes to submit proposals for more than one specific height to every primary block, and then nodes would pick the first proposal for each height. Thus, allowing nodes to log proposals at a higher but constant rate. This may satisfy progress, but doesn’t satisfy No-Batching Validity, as nodes have to propose several blocks at once, and proposals may contain transactions which depend on previous ones to be accepted.
Another option is to switch committees at set-length epochs, say, 1000 blocks (assuming $\Delta_{\textit{active}}$ is large enough). Before the end of an epoch, the committee writes back the last value, and the next committee picks up from that point. Again, this does not fulfill Progress property, since during the handoff the chain must stop for about $O(\Delta_{\underline{\mathcal{L}}\textit{-write}})$ until the previous-epoch value is written.
4.2. Impossibility Proof
First, before we show the impossibility, we show any solution which satisfies No-Batching Validity must allow no batching.
Lemma 1.
Let $\mathcal{A}$ be a protocol, and let $\sigma$ be an execution of $\mathcal{A}$ where all nodes are correct. Let $t_{k}$ be the time when the first node decides on height $k$, and let $t_{k+1}$ be the time when the first node decides on height $k+1$. If $\mathcal{A}$ satisfies No-Batching Validity, then $t_{k+1}>t_{k}$.
Proof.
Let $\mathcal{A}$ be a protocol that satisfies No-Batching Validity, and let $\sigma$ be an execution of the protocol where all nodes are correct. Let $t_{k}$ be the time when the first node decides on height $k$, and let $t_{k+1}$ be the time when the first node decides on height $k+1$. Assume the value decided for height $k+1$ is $v$.
Assume towards contradiction that $t_{k+1}\leq t_{k}$. Consider an execution $\sigma^{\prime}$ identical to $\sigma$ up to $t_{k+1}$, and afterward, all input functions for height $k+1$ and times ${t^{\prime\prime}>t_{k}}$ return ${v^{\prime}\neq v}$, all nodes are correct, and there are no configuration changes. By No-Batching Validity, no node decides on a value different from $v^{\prime}$ at height $k+1$. But $v\neq v^{\prime}$ by definition, so this contradicts the assumption that $t_{k+1}\leq t_{k}$. ∎
We are now ready to show that no solution is possible without synchronous access to the primary ledger.
Theorem 2.
No protocol implements a tethered chain with partially synchronous ledger writing, when facing Byzantine behavior of inactive (unstaked) nodes.
Proof.
Assume towards contradiction that there is a protocol $\mathcal{A}$ satisfying Agreement, No-Batching Validity, and Progress.
The proof uses a set of executions and indistinguishability arguments. Figure 2 illustrates the executions.
Let $\sigma$ be an execution of $\mathcal{A}$ where all nodes are correct. Let $t_{\textit{GST}}$ be the time from which network communication becomes stable in the execution $\sigma$. As in the Progress property, the protocol starts progressing rapidly after time $t_{\textit{GST}}+\Delta_{\textit{stabilize}}$. Denote by $k$ the largest height decided by any node by that time. Define $t^{\textit{divergence}}$ as the earliest time any node logs height $k+1$. Since no node can log height $k+1$ before completing height $k$ (Lemma 1), we have $t^{\textit{divergence}}>t_{\textit{GST}}+\Delta_{\textit{stabilize}}$.
We now define two executions, $\sigma^{0}$ and $\sigma^{1}$, which follow $\sigma$ exactly until $t^{\textit{divergence}}$ but diverge afterward. In $\sigma^{0}$, all input functions for height $k+2$ return 0 after $t^{\textit{divergence}}$. In $\sigma^{1}$, they all return 1 after $t^{\textit{divergence}}$. We assume that no committee changes occur after $t^{\textit{divergence}}$ in either execution, that all nodes remain correct and that both executions $t_{\textit{GST}}$ is the same as in $\sigma$. We also assume that the scheduler delivers all node-to-node messages within a single step and completes primary writes at the latest possible time, within $\Delta_{\underline{\mathcal{L}}\textit{-write}}$ steps of each write. By Progress, in both executions $\sigma^{0}$ and $\sigma^{1}$, at least one node decides on height $k+2$ by time $t^{\textit{decide}}=t^{\textit{divergence}}+\Delta_{\textit{consensus}}+\Delta_{\textit{prop}}$. By No-Batching Validity, in $\sigma^{0}$ one node logs 0, while in $\sigma^{1}$ one node (not necessarily the same one) logs 1.
Any write to the primary ledger between $t^{\textit{divergence}}$ and $t^{\textit{decide}}$ can take up to $\Delta_{\underline{\mathcal{L}}\textit{-write}}$ steps, and we assume $\Delta_{\underline{\mathcal{L}}\textit{-write}}$ is large enough compared to $\Delta_{\textit{consensus}}+\Delta_{\textit{prop}}$ so that none of these writes occur by time $t^{\textit{decide}}$.
Next, define a third execution, $\sigma^{1^{\prime}}$, which follows $\sigma^{1}$ until $t^{\textit{decide}}$ and then diverges. All nodes in the old committee issue unstake transactions at time ${t^{\textit{decide}}}$ and a disjoint set of nodes stakes in to form a new committee at time ${t^{\textit{decide}}+\Delta_{\textit{active}}}$. The scheduler delivers the unstaking and staking transactions as soon as possible. Set $t_{\textit{GST}}^{\prime}=t^{\textit{decide}}+\Delta_{\textit{active}}$ for the execution $\sigma^{1^{\prime}}$. The scheduler delays all messages from the old committee to the new committee and all message writes until the latest possible time $t_{\textit{GST}}^{\prime}+\Delta_{\textit{prop}}$, and $t_{\textit{GST}}^{\prime}+\Delta_{\underline{\mathcal{L}}\textit{-write}}$ respectively. By Progress, the new committee, now active, logs a value at height $k+2$ by some time $t^{\textit{decide}^{\prime}}>t^{\textit{decide}}+\Delta_{\textit{active}}$; by Agreement, that value is 1.
Finally, we construct $\sigma^{x}$ by combining the prefix of $\sigma^{0}$ up to $t^{\textit{decide}}$ with the reconfiguration at time $T$ that occurs in $\sigma^{1^{\prime}}$. In $\sigma^{x}$, the old committee has already logged 0 for height $k+2$ by $t^{0}$. We set the stabilization time of $\sigma^{x}$ to $t_{\textit{GST}}^{x}=t^{\textit{decide}^{\prime}}$, and let the scheduler defer all old-committee messages and ledger writes which occur after $t^{\textit{decide}}$ until $t_{\textit{GST}}^{x}$, except for the following. Since by $T+\Delta_{\textit{active}}$ ($<t_{\textit{GST}}^{x}$) the old committee is no longer active, it can behave arbitrarily without being slashed. At that point, it produces the same messages and write orders as it did in $\sigma^{1^{\prime}}$, which the scheduler delivers at the corresponding times they appeared in $\sigma^{1^{\prime}}$. Hence, the new committee in $\sigma^{x}$ sees exactly the events up until $t^{\textit{decide}^{\prime}}$ as in $\sigma^{1^{\prime}}$, so it cannot distinguish $\sigma^{x}$ from $\sigma^{1^{\prime}}$ and therefore logs 1 at height $k+2$ at time $t^{\textit{decide}^{\prime}}$. This contradicts Agreement, because the same height $k+2$ is logged as both 0 and 1 in a single execution ($\sigma^{x}$).
Therefore, no protocol can achieve a tethered chain under partially synchronous primary chain writing when inactive nodes may behave arbitrarily. ∎
[…]
Illustration of impossibility proof.
Figure 2. Impossibility proof.
5. Aegis
@@
After establishing that no secure tethered chain can be achieved in a partially synchronous setting, we present Aegis, a tethered-chain protocol that takes advantage of synchronous access to a primary ledger. We begin with an overview of Aegis (§5.1). Then we describe its components in detail: Aegis’s blockchain structure (§5.2), its primary-chain contract (§5.3), and the peer-to-peer protocol that Aegis nodes execute to form the blockchain (§5.4). Finally, we discuss the execution of the forensics procedure to identify and penalize misbehaving nodes (§5.5).
5.1. Protocol Overview
Aegis Blockchain.
The Aegis blockchain advances by blocks generated by committees. Committees use a Byzantine fault-tolerant consensus protocol designed for partial synchrony. Figure 3 illustrates Aegis’s blockchain topology and its key mechanisms. Each Aegis block contains hash references to its parent block and to a primary-chain block. The parent reference establishes the chain’s history, while the primary-chain reference defines the committee responsible for producing the next block. For instance, Aegis block $A1$ references primary block $P2$, so $P2$’s state determines the committee for block $A2$. Similarly, $A2$’s reference to $P2$ means $P2$’s state determines the committee for block $A3$.
[…]
Figure 3. Aegis’s toplogy and key mechanisms.
Committee Selection.
A primary block defines a committee from nodes with locked stake weighted by their stake amount.
Note 1.
Committees can be selected in various ways to utilize the consensus protocol more efficiently. For example, committees can include only nodes exceeding a minimum stake threshold or a fixed number of highest-stake nodes.
Note 2.
In staking transactions, nodes must include contact information (e.g., IP address) for message routing; we omit this detail for simplicity.
Aegis handles committee transitions without explicit handovers through references to primary blocks. As shown in Figure 3, committee 1 initializes committee 2 as block $A2$ references primary block $P2$. Each committee remains active for a bounded period ($\Delta_{\textit{active}}$) after the referenced primary block. During this period, members cannot withdraw their stake. This provides the accountability necessary for safety, meaning that the chain is agreed upon by all correct nodes and cannot be altered retroactively.
Primary-Chain Smart Contract.
In addition to primary block references, Aegis employs two complementary mechanisms through its primary-chain smart contract: checkpoints and resets.
The checkpoint mechanism anchors the Aegis state to the primary chain, creating tamper-proof reference points that (1) help new nodes determine the correct chain history and (2) signal the smart contract that the Aegis chain is currently progressing. Active committees submit a checkpoint before the end of their activity period. For instance, a node in committee 1 submits a checkpoint of block $A3$ to $P3$. The smart contract verifies that the block’s committee is active during submission, that the block results from valid consensus, and that it is a descendant of the last checkpointed block.
The reset mechanism helps Aegis recover from communication failures. If network delays prevent a committee from producing blocks before its activity period expires, no new committee will be selected. In this case, any node can issue a reset through the primary-chain contract. A reset establishes a new committee based on the current primary-chain state, which can then resume extending the Aegis chain from the last checkpointed block. The committee produces a block that includes a hash reference to the primary-chain block containing the reset. For example, after committee 2’s activity period ended, a reset was submitted to block $P4$, creating a committee that produced a block referencing $P4$.
Node Protocol.
Aegis nodes execute a distributed protocol that utilizes the primary chain, the smart contract, and its checkpoint and reset mechanisms. Each node maintains a local view of the Aegis blockchain, coordinated with other nodes via peer-to-peer communication.
Nodes follow the latest state of the primary chain: they reference the latest block when producing an Aegis block and monitor the smart contract for checkpoints and resets. Nodes have access to previous primary-chain blocks when needed and can post checkpoints and resets to the contract. Aegis relies on these actions taking at most $\Delta_{\underline{\mathcal{L}}\textit{-write}}$ time steps.
The protocol operates in three sequential phases at each time step. First, nodes read the primary chain to identify the latest checkpointed block and determine the active committee. Then, they log all Aegis blocks created by active committees since the last checkpoint. Finally, nodes run a Byzantine fault tolerant consensus protocol to extend the Aegis chain.
Timing Constraints.
There are two challenges that might arise when a committee approaches the end of its activity period. First, a committee might benignly fail to checkpoint a block or forward it to the next committee. This would result in the next committee creating a conflicting block that does not reference the previous committee’s block, violating agreement. Second, committee nodes might maliciously produce conflicting blocks right before their period ends, leaving insufficient time for forensics and penalties, removing incentives to behave properly.
Aegis addresses these challenges with strict timing constraints to ensure sufficient time for checkpoints and forensics. Near the end of the activity period, nodes stop producing new blocks. Instead, they checkpoint their last generated block to ensure proper handover, then wait for their tenure to end. This period covers the last $3\Delta_{\underline{\mathcal{L}}\textit{-write}}$ steps of each committee’s activity period.
Another challenge is premature resets that could create concurrent committees and conflicting blocks. To prevent this, the smart contract accepts resets only when the latest checkpoint or reset is at least $\Delta_{\textit{active}}$ time steps old. For example, since $P1$ contains a reset, no other reset is accepted during committee 1’s activity period, lasting $\Delta_{\textit{active}}$ steps from $P1$’s creation. Additionally, a checkpoint to $P3$ restarts this restriction for $\Delta_{\textit{active}}$ steps following the checkpoint.
Next, we describe the components of Aegis in detail, starting with the data structure of the Aegis blockchain.
5.2. Chain Structure